# A Survey for Deep Reinforcement Learning Based Network Intrusion Detection

### v1

@@ IV-D Improved network feature engineering
Typically, the dataset of network traffic comprises various features that reflect the status of network traffic. However, not all of these features are beneficial for constructing DRL-based network intrusion detection systems. The accurate representation of traffic status using network traffic features is crucial. Consequently, many researches have been proposed to effectively extract network features that can accurately represent the actual status of network traffic. Liu et al [95] proposed a method that combines Local-Sensitive Hashing (LSH) with Deep Convolutional Neural Networks (DCNN), selecting optimal features by assessing the distribution of information entropy across each feature value. Ren et al [93] suggest combining the Recursive Feature Elimination (RFE) and decision tree to select the optimal sub-feature set and the method effectively identifies and eliminates approximately 80% of the redundant features from the original dataset. Ren et al [81] further employed Graph Convolutional Networks (GCN) to extract deep features from network data. They transformed the selected input data into dynamic graph networks. Through the hierarchical structure of GCN, more rich and abstract features were extracted. Finally, they combined a multi-agent learning framework to transform the traditional feature selection space of $2^{N}$ into a competition among $N$ feature agents, effectively reducing the feature space.
IV-E Handling unbalanced datasets
@@
The volume of normal network traffic data significantly exceeds that of intrusion data, which is a common-sense observation. Consequently, nearly all network intrusion detection datasets suffer from a severe imbalance in attack-type distribution. Therefore, the challenge of training effective DRL-based NID models on an imbalanced dataset have consistently attracted attention. Researchers have attempted to propose various methods to address such issues. Lopez et al [84] addressed this problem in NID by augmenting Radial Basis Function (RBF) [96] neural networks and integrating them with offline reinforcement learning algorithms. They validated the superior performance of this approach across five commonly used datasets. However, the proposed method may face challenges with larger action spaces and have higher computational costs in training. Mohamed et al [97] utilized a deep State-Action-Reward-State-Action (SARSA) algorithm [98] combined with Deep Neural Networks (DNN) [99] to address the issue of data imbalance in NIDs. Although the performance was outstanding, the authors did not analyze the potential drawbacks of their proposed SARSA algorithm. Caution should be exercised when applying this approach. Pashaei et al [94] introduced an adversarial DRL model combined with intelligent environment simulation, presenting an effective approach to addressing the issue of high-dimensional data imbalance in NIDs. This method improves overall classification performance by increasing attention to minority classes, particularly demonstrating its efficiency and effectiveness in practical applications within IoT environments.
@@ IV-F Training efficiency
Normally, network traffic data exhibit high levels of uncertainty and complexity, leading to low training efficiency of DRL, which has been a long-standing concern. Louati et al [79] developed a distributed multi-agent reinforcement learning approach for distributed intrusion detection in large-scale network environments, termed Big-IDS. While the model demonstrated impressive performance, a notable drawback is its low training efficiency. Training the model takes approximately three days when encryption is used and about 12 hours without encryption. This significant time requirement highlights the need for optimization in training procedures to enhance practical applicability. Li et al [80] introduced a network intrusion detection model named AE-SAC, based on adversarial environment learning and the Soft Actor-Critic (SAC) DRL algorithm. While AE-SAC achieved excellent performance in terms of accuracy and F1 score, its complex network architecture resulted in extended training time. During each training session, both the environment agent and the classifier agent are required to update at least three networks, contributing to the lengthy training process. Kalinin et al [100] enhanced the training efficiency of deep reinforcement learning models in Internet of Things(IoT) intrusion detection by implementing lightweight neural network architectures and developing various multi-agent system architectures. This approach also demonstrated superior performance in terms of accuracy and completeness metrics.
IV-G Identifying minority and unknown attacks
@@
Normal network traffic constitutes the major class of the these mentioned dataset. However, in practical applications, the robustness and generalization ability of NID systems are often of greater concern. Therefore, identifying minority types of attacks and recognizing unknown categories of attacks are crucial in the actual deployment of NID systems. Hsu et al [90]developed a DRL-based model for network intrusion detection, equipped with detection and learning modes. This model can switch flexibly based on network traffic behaviours, enabling self-updating capabilities that enhance its ability to recognize unknown network traffic. The study was also tested in a real network environment, where it demonstrated good performance as well. Malika et al [101] introduced a distributed multi-agent NID system that combines DRL with attention mechanisms [102]. Utilizing Distributed Q-networks (DQNs) deployed across multiple network nodes, this system offers varied perspectives on the network’s security status. It features a multi-agent attention mechanism that increasingly focuses on specific network nodes as performance improves. Additionally, the design incorporates zero-day defence measures to mitigate attacks exploiting unknown vulnerabilities. Notably, the integration of denoising autoencoders (DAE) has significantly enhanced the model’s performance. Liu et al [89] enhanced a DRL framework by incorporating human operator interaction feedback into the MDP, creating a hybrid structure of Q-networks. They integrated Long Short-Term Memory (LSTM) [103] networks to better handle time-series features and incorporated a prioritized experience replay mechanism. This innovative approach significantly improved the recognition capabilities for minority and unknown category attacks. Another rising issue is the trade-off between identifying minority attacks and unknown attacks. Dong et al [85] have reported that when they enhancing the capability to recognize unknown attacks, they have unfortunately seen a decline in the ability to identify specific minority attack types. Xiangyu et al [87] developed a novel method that combines an enhanced version of the Synthetic Minority Over-sampling Technique (SMOTE) with adversarial RL to improve the detection accuracy of minority classes, such as anomalies or types of attacks. By using SMOTE to generate synthetic samples, they increased the representation of minority classes, thereby addressing the issue of dataset imbalance. Furthermore, the use of two agents within the RL framework - a classifier agent and an environment agent to facilitate dynamic training and data selection, further optimizing the model’s ability to recognize minority classes.
@@ V-B Models
Despite the initial exploration of DRL in NID that has been implemented recently, it is noteworthy that the structural composition of most DRL policies are still based on simple MLP. According to current survey [104], even the most popular transformer architecture has not been widely introduced in DRL-based NID systems, with only a few models like [101] employing attention mechanisms and achieving good results. One possible reason is that traditional supervised learning and deep learning approaches continue to receive more attention from the network intrusion detection community [105, 106]. Additionally, current NID datasets intuitively seem more suited for supervised learning methods [107]. Recently, the emergence of a novel architecture called Kolmogorov–Arnold Networks(KAN) [108], inspired by Kolmogorov-Arnold representation theorem [109], has attracted widespread attention in general AI community. This architecture shifts activation functions traditionally located on MLP neurons to the connections/edges between neurons. In their experiments, it replicated the results of a 300,000-parameter MLP-based neural network model using fewer than 200 parameters. Thus, in the foreseeable future, whether KAN could be used to enhance DRL-based NID systems is an interesting topic, especially since most DRL-based NID systems currently approximate optimal strategies using simple MLP structures. And DRL-based NID seems lagging behind the latest DRL models or methods. Advanced techniques such as PPO [43] and TRPO [42] have not been embraced by the NID community. DRL-based NID is still in its early stages with many classic DRL-based algorithms. Lastly, the evaluation of the generalization of DRL-based NID models are limited, and relying solely on training and testing splits from a single dataset appears insufficient. In our survey, only a few studies not only trained models but also conducted experiments in real network scenarios [90]. Real-world deployments and tests are urgently needed for DRL-based NID.
V-C Inverse reinforcement learning for NID
@@
Traditional DRL is known for its inefficiency in sample utilization [110, 111]. The advent of inverse reinforcement learning may have potential to expedite policy optimization for agents [112], it is also critically important for NID systems that require high level of real-time responsiveness. However, current research indicates that IRL has not been applied within the context of DRL-based NIDS. Only some related attempts like Liu et al [113] successfully used Inverse Reinforcement Learning (IRL) to reverse-engineer the reward function by observing the trajectories of a trained DQN controller and used this function to design attack strategies that effectively disrupt the control functions of an Industrial Internet of Things (IIoT) system. One future possible way could involve treating existing network intrusion datasets as expert demonstration data and employing IRL methodologies to model the underlying reward functions of various attacker behaviours. Subsequently, these specialized reward functions could be utilized to train defender agents. For datasets collected from actual network attacks, IRL can be used to effectively model the behaviours of real attackers and potentially train more robust defender agent. Despite this promising research direction, the inherent computational complexity of IRL [50] algorithms may need the introduction of specific methods for improvement to actualize the theoretical concepts proposed.
V-D DRL for intrusion detection to IoT
@@
In recent years, there has been more and more research focusing on the network detection forIoT [114, 115]. Haosen et al had made an initial brief survey on the DRL-based intrusion detection in IoT[116]. And, as the number of smart devices increases, there is growing concern about how to prevent network intrusions targeting IoT devices[117]. Current approaches are predominantly based on traditional deep learning and machine learning methods as well[118]. There is a notable lack of exploration in IoT-specific DRL-based techniques. And in our survey, only [100, 119], [87] and [113] have focused on intrusion detection specifically for IoT devices using DRL/RL in recent years. One possible reason for this is the scarcity of benchmark datasets suitable for IoT and DRL. To date, researchers have only conducted preliminary studies on datasets like AWID [76], Bot-IoT[120], and IoTID20 [121]. Nevertheless, with the explosive growth of IoT devices [122], the demand for reliable intrusion detection technologies continues to rise. Therefore, there should remain an optimistic outlook on the application of DRL techniques in IoT intrusion detection. Additionally, researchers from the broad IoT intrusion detection community should continue to strive towards proposing new datasets specifically tailored for IoT device intrusion detection. Moreover, some surveys have also revealed several challenges when introducing DRL into IoT applications, including the need to handle continuous state-action spaces, learning in partially observable environments, ensuring robustness against adversarial attacks, real-time decision making and data privacy and security [123, 124].
V-E Generative deep model enhanced DRL for NID
@@
Generative artificial intelligence, such as large language models (LLMs) [125], has recently garnered significant attention worldwide. While some researchers have noted the transformative potential of these methods in the broad field of cybersecurity, few studies have explored how generative AI can be integrated with DRL to advance research in NID. Some researchers have already recognized that LLMs can enhance DRL models in areas such as reward design and world model simulator [126]. Consequently, in the foreseeable future, employing LLMs and other generative AI models to augment RL approaches is likely to become a new research paradigm. Therefore, exploring how this potential research paradigm could be extended to NID may become an intriguing research hotspot. Combining the powerful capability of LLM-based world models [127, 128], One initial idea is to fine-tuning LLM as a network intrusion expert (attacker)[129, 130] and then using the principle of DRL to train the defender agent though the interaction with LLM-based attacker. Then, using the current public dataset to test and make an evaluation of the defenders. Additionally, the widespread availability of generative AI might also inspire traditional attackers to change their way of attack [130]. For instance, with the aid of LLMs, attackers could potentially improve and optimize their attack strategies [130], significantly degrading the performance of NID systems trained on conventional datasets, an urgent concern that should be addressed by the community. Moreover, we cannot guarantee that these potential attackers will cease using these tools as AI technology advances. However, one thing that should be clear is researchers in NID defence systems should responsibly learn to use generative AI to enhance the reliability and defensive capabilities of their proposed systems and make reliable tests. Moreover, due to large simulated datasets and fewer real attack examples exist. LLMs attacker might be, to some extend, fitting only to the simulated scenario if the LLMs lack enough knowledge and logic transferring and generalizing ability. It should be a remaining concern in the coming future anyway.
V-F Policy functions and architectures
@@
The findings over recent years suggest that deep neural networks can indeed represent complex policy functions within DRL [131]. Deep learning has made significant advancements, greatly enhancing agents’ capabilities in learning and generalizing from complex environments, which is particularly relevant to network security. Previous surveys also identified a growing interest in integrating attention mechanisms into policy function representations. The promise of attention mechanisms lies in their ability to handle dynamic and high-dimensional data, which is a typical feature in network traffic analysis. However, it’s important to note that research on deep model architectures is progressing rapidly. For example, architectures like Mamba [132] and Learn at Test Time (TTT) [133] structures have gained popularity recently due to their innovative design and superior performance metrics. These architectures also offer increased capabilities in model interpretability, scalability, and robustness, which are critical for developing effective intrusion detection systems. The need to address this through advanced policy function representations should be a high priority for the network intrusion detection community. By embedding these state-of-the-art deep learning architectures into IDS, researchers may overcome the limitations of conventional methods, offering more precise and efficient threat identification. Implementing such architectures could result in a new generation of IDS that are far better equipped to protect against today’s sophisticated cyber threats in our increasingly networked and complex digital environment.

### v2

@@ IV-E Improved network feature engineering
Typically, the dataset of network traffic comprises various features that reflect the status of network traffic. However, not all of these features are beneficial for constructing DRL-based network intrusion detection systems. The accurate representation of traffic status using network traffic features is crucial. Consequently, many research efforts have been proposed to effectively extract network features that can accurately represent the actual status of network traffic. Liu et al. [63] proposed a method that combines Local-Sensitive Hashing (LSH) with Deep Convolutional Neural Networks (DCNN), selecting optimal features by assessing the distribution of information entropy across each feature value. Ren et al. [100] suggest combining the Recursive Feature Elimination (RFE) and decision tree to select the optimal sub-feature set and the method effectively identifies and eliminates approximately 80% of the redundant features from the original dataset. Ren et al. [101] further employed Graph Convolutional Networks (GCN) to extract deep features from network data. They transformed the selected input data into dynamic graph networks. Through the hierarchical structure of GCN, more rich and abstract features were extracted. Finally, they combined a multi-agent learning framework to transform the traditional feature selection space of $2^{N}$ into a competition among $N$ feature agents, effectively reducing the feature space.
IV-F Handling unbalanced datasets
@@
The volume of normal network traffic data significantly exceeds that of intrusion data, which is a common-sense observation [33]. Consequently, most network intrusion detection datasets suffer from a severe imbalance in attack-type distribution [1]. Therefore, the challenge of training effective DRL-based NID models on an imbalanced dataset have consistently attracted attention. Researchers have attempted to propose various methods to address such issues. Lopez et al. [66] addressed this problem in NID by augmenting Radial Basis Function (RBF) [11] neural networks and integrating them with offline reinforcement learning algorithms. They validated the superior performance of this approach across five commonly used datasets. However, the proposed method may face challenges with larger action spaces and have higher computational costs in training. Mohamed et al. [77] utilized a deep State-Action-Reward-State-Action (SARSA) algorithm [104] combined with Deep Neural Networks (DNN) [9] to address the issue of data imbalance in NIDs. Although the performance was outstanding, the authors did not analyze the potential drawbacks of their proposed SARSA algorithm. Caution should be exercised when applying this approach. Pashaei et al. [96] introduced an adversarial DRL model combined with intelligent environment simulation, presenting an effective approach to addressing the issue of high-dimensional data imbalance in NIDs. This method improves overall classification performance by increasing attention to minority classes, particularly demonstrating its efficiency and effectiveness in practical applications within IoT environments.
@@ IV-G Training efficiency
Normally, network traffic data exhibit high levels of uncertainty and complexity [60], leading to low training efficiency of DRL, which has been a long-standing concern [83]. Louati et al. [67] developed a distributed multi-agent reinforcement learning approach for distributed intrusion detection in large-scale network environments, termed Big-IDS. While the model demonstrated impressive performance, a notable drawback is its low training efficiency. Training the model takes approximately three days when encryption is used and about 12 hours without encryption. This significant time requirement highlights the need for optimization in training procedures to enhance practical applicability. Li et al. [59] introduced a network intrusion detection model named AE-SAC, based on adversarial environment learning and the Soft Actor-Critic (SAC) DRL algorithm. While AE-SAC achieved excellent performance in terms of accuracy and F1 score, its complex network architecture resulted in extended training time. During each training session, both the environment agent and the classifier agent are required to update at least three networks, contributing to the lengthy training process. Kalinin et al. [47] enhanced the training efficiency of deep reinforcement learning models in Internet of Things (IoT) intrusion detection by implementing lightweight neural network architectures and developing various multi-agent system architectures. This approach also demonstrated superior performance in terms of accuracy and completeness metrics.
IV-H Identifying minority and unknown attacks
@@
Normal network traffic constitutes the major class of the these mentioned dataset. However, in practical applications, the robustness and generalization ability of NID systems are often of greater concern. Therefore, identifying minority types of attacks and recognizing unknown categories of attacks are crucial in the actual deployment of NID systems. Hsu et al. [39]developed a DRL-based model for network intrusion detection, equipped with detection and learning modes. This model can switch flexibly based on network traffic behaviours, enabling self-updating capabilities that enhance its ability to recognize unknown network traffic. The study was also tested in a real network environment, where it demonstrated good performance as well. Malik et al. [71] introduced a distributed multi-agent NID system that combines DRL with attention mechanisms [135]. Utilizing Distributed Q-networks (DQNs) deployed across multiple network nodes, this system offers varied perspectives on the network’s security status. It features a multi-agent attention mechanism that increasingly focuses on specific network nodes as performance improves. Additionally, the design incorporates zero-day defence measures to mitigate attacks exploiting unknown vulnerabilities. Notably, the integration of denoising autoencoders (DAE) has significantly enhanced the model’s performance. Liu et al. [62] enhanced a DRL framework by incorporating human operator interaction feedback into the MDP, creating a hybrid structure of Q-networks. They integrated Long Short-Term Memory (LSTM) [31] networks to better handle time-series features and incorporated a prioritized experience replay mechanism. This innovative approach significantly improved the recognition capabilities for minority and unknown category attacks. Another rising issue is the trade-off between identifying minority attacks and unknown attacks. Dong et al. [24] have reported that when they enhancing the capability to recognize unknown attacks, they have unfortunately seen a decline in the ability to identify specific minority attack types. Xiangyu et al. [69] developed a novel method that combines an enhanced version of the Synthetic Minority Over-sampling Technique (SMOTE) with adversarial RL to improve the detection accuracy of minority classes, such as anomalies or types of attacks. By using SMOTE to generate synthetic samples, they increased the representation of minority classes, thereby addressing the issue of dataset imbalance. Furthermore, they employed two agents within the RL framework, a classifier agent and an environment agent, to facilitate dynamic training and data selection, further optimizing the model’s ability to recognize minority classes.
@@ V-B Models
Despite recent initial explorations of DRL in NID, it is noteworthy that the structural composition of most DRL policies are still based on simple MLP. According to a recent survey [2], even the most popular transformer architecture has not been widely introduced in DRL-based NID systems, with only a few models like [71] employing attention mechanisms and achieving good results. One possible reason is that traditional supervised learning and deep learning approaches continue to receive more attention from the network intrusion detection community [144, 37]. Additionally, current NID datasets intuitively seem more suited for supervised learning methods [127]. Recently, the emergence of a novel architecture called Kolmogorov–Arnold Networks (KAN) [64], inspired by Kolmogorov-Arnold representation theorem [52], has attracted widespread attention in general AI community. This architecture shifts activation functions traditionally located on MLP neurons to the connections/edges between neurons. In their experiments, it replicated the results of a 300,000-parameter MLP-based neural network model using fewer than 200 parameters. Thus, in the foreseeable future, whether KAN could be used to enhance DRL-based NID systems is an interesting topic, especially since most DRL-based NID systems currently approximate optimal strategies using simple MLP structures. Moreover, DRL-based NID seems to lag behind the latest DRL models or methods. Advanced techniques such as PPO [108] and TRPO [107] have not been embraced by the NID community. DRL-based NID is still in its early stages with many classic DRL-based algorithms. Lastly, the evaluation of the generalization of DRL-based NID models are limited, and relying solely on training and testing splits from a single dataset appears insufficient. In our survey, only a few studies not only trained models but also conducted experiments in real network scenarios [39]. Real-world deployments and tests are urgently needed for DRL-based NID.
V-C Tools and Simulation Environments
Table I introduced various reinforcement learning tools and simulation environments that exhibit significant heterogeneity in their purposes and applicability to network intrusion detection (NID) tasks. A notable example is YAWNING-TITAN, which is primarily designed for Autonomous Cyber Defence rather than intrusion detection, focusing on active defence operations such as node isolation and vulnerability patching—a fundamentally different problem space from the passive traffic monitoring and anomaly detection that characterizes traditional NID. Similarly, several other tools are oriented toward penetration testing and offensive security operations rather than intrusion detection. This heterogeneity presents both challenges and opportunities for researchers. While careful selection of simulation environments is essential to ensure alignment with NID requirements, the diverse ecosystem also enables cross-domain learning, where adversarial strategies from penetration testing environments could inform more robust NID systems, and defensive tactics from autonomous cyber defence could inspire novel active learning approaches. Moving forward, the community would benefit from establishing clearer taxonomies that distinguish tools by their primary objectives—detection, defence, or offence—and developing simulation environments specifically tailored to NID tasks with realistic network traffic generation and evaluation metrics aligned with detection performance.
V-D Inverse reinforcement learning for NID
@@
Traditional DRL is known for its inefficiency in sample utilization [146, 148]. The advent of inverse reinforcement learning may have potential to expedite policy optimization for agents [21], it is also critically important for NID systems that require high level of real-time responsiveness. However, current research indicates that IRL has not been applied within the context of DRL-based NIDS. Only some related attempts like Liu et al. [61] successfully used Inverse Reinforcement Learning (IRL) to reverse-engineer the reward function by observing the trajectories of a trained DQN controller and used this function to design attack strategies that effectively disrupt the control functions of an Industrial Internet of Things (IIoT) system. One possible future direction could involve treating existing network intrusion datasets as expert demonstration data and employing IRL methodologies to model the underlying reward functions of various attacker behaviours. Subsequently, these specialized reward functions could be utilized to train defender agents. For datasets collected from actual network attacks, IRL can be used to effectively model the behaviours of real attackers and potentially train more robust defender agent. Despite this promising research direction, the inherent computational complexity of IRL [142] algorithms may need the introduction of specific methods for improvement to actualize the theoretical concepts proposed.
V-E DRL for intrusion detection to IoT
@@
In recent years, there has been an increase in research focusing on the network detection for IoT [73, 105]. Haosen et al. made an initial brief survey on the DRL-based intrusion detection in IoT[147]. As the number of smart devices increases, there is growing concern about how to prevent network intrusions targeting IoT devices[75]. Current approaches are predominantly based on traditional deep learning and machine learning methods as well[50]. There is a notable lack of exploration in IoT-specific DRL-based techniques. In our survey, only [47, 38], [69] and [61] have focused on intrusion detection specifically for IoT devices using DRL/RL in recent years. One possible reason for this is the scarcity of benchmark datasets suitable for IoT and DRL. To date, researchers have only conducted preliminary studies on datasets like AWID [15], Bot-IoT[95], and IoTID20 [132]. Nevertheless, with the explosive growth of IoT devices [120], the demand for reliable intrusion detection technologies continues to rise. Therefore, there should remain an optimistic outlook on the application of DRL techniques in IoT intrusion detection. Additionally, researchers from the broad IoT intrusion detection community should continue to strive towards proposing new datasets specifically tailored for IoT device intrusion detection. Moreover, some surveys have also revealed several challenges when introducing DRL into IoT applications, including the need to handle continuous state-action spaces, learning in partially observable environments, ensuring robustness against adversarial attacks, real-time decision making and data privacy and security [16, 133].
V-F Generative deep model enhanced DRL for NID
@@
Generative artificial intelligence, such as large language models (LLMs) [14], has recently garnered significant attention worldwide. While some researchers have noted the transformative potential of these methods in the broad field of cybersecurity, few studies have explored how generative AI can be integrated with DRL to advance research in NID. Some researchers have already recognized that LLMs can enhance DRL models in areas such as reward design and world model simulator [13]. Consequently, in the foreseeable future, employing LLMs and other generative AI models to augment RL approaches is likely to become a new research paradigm. Therefore, exploring how this potential research paradigm could be extended to NID may become an intriguing research hotspot. Given the powerful capability of LLM-based world models [29, 145], one initial idea is to fine-tuning a LLM as a network intrusion expert (attacker)[82, 26] and then using the principle of DRL to train the defender agent through the interaction with LLM-based attacker. Then, using the current public dataset to test and make an evaluation of the defenders. Additionally, the widespread availability of generative AI might also inspire traditional attackers to change their way of attack [26]. For instance, with the aid of LLMs, attackers could potentially improve and optimize their attack strategies [26], significantly degrading the performance of NID systems trained on conventional datasets, an urgent concern that should be addressed by the community. Moreover, we cannot guarantee that these potential attackers will cease using these tools as AI technology advances. However, one thing that should be clear is researchers in NID defence systems should responsibly learn to use generative AI to enhance the reliability and defensive capabilities of their proposed systems and make reliable tests as well. Moreover, due to the fact that large simulated datasets exist but fewer real attack examples are available. LLMs attacker might be, to some extent, fitting only to the simulated scenario if the LLMs lack enough knowledge and logic transferring and generalizing ability. It should be a remaining concern in the coming future anyway.
V-G Policy functions and architectures
@@
The findings over recent years suggest that deep neural networks can indeed represent complex policy functions within DRL [57]. Deep learning has made significant advancements, greatly enhancing agents’ capabilities in learning and generalizing from complex environments, which is particularly relevant to network security. Previous surveys also identified a growing interest in integrating attention mechanisms into policy function representations. The promise of attention mechanisms lies in their ability to handle dynamic and high-dimensional data, which is a typical feature in network traffic analysis. However, it is important to note that research on deep model architectures is progressing rapidly. For example, architectures like Mamba [32] and Learn at Test Time (TTT) [125] structures have gained popularity recently due to their innovative design and superior performance metrics. These architectures also offer increased capabilities in model interpretability, scalability, and robustness, which are critical for developing effective intrusion detection systems. The need to address this through advanced policy function representations should be a high priority for the network intrusion detection community. By embedding these state-of-the-art deep learning architectures into IDS, researchers may overcome the limitations of conventional methods, offering more precise and efficient threat identification. Implementing such architectures could result in a new generation of IDS that are far better equipped to protect against today’s sophisticated cyber threats in our increasingly networked and complex digital environment.
V-H Risks of autonomous AI agents in network intrusion
It is worth noting that with the rapid advancement of autonomous AI agent technologies, the risk of their malicious exploitation for automated network intrusion attacks is becoming increasingly prominent. On one hand, autonomous agents trained via deep reinforcement learning have already demonstrated the capability to automatically probe vulnerabilities and plan attack paths in penetration testing scenarios [40]. On the other hand, LLM-based open-source autonomous agents, exemplified by OpenClaw [123], have recently attracted widespread attention. Such agents are capable of autonomously executing system commands, browser automation, and file management operations, with their open-source nature and extensible plugin ecosystems providing remarkable functional flexibility. However, security researchers have already identified critical vulnerabilities in such agent frameworks, including remote code execution (e.g., CVE-2026-25253 [84]) and data exfiltration and prompt injection risks within the skill supply chain. When such autonomous agents—whether driven by deep reinforcement learning or large language models—are deliberately leveraged by adversaries, they could autonomously identify target network vulnerabilities, formulate attack strategies, and execute intrusion operations without human intervention, posing a serious challenge to existing network intrusion detection systems [26]. Therefore, while the cybersecurity community actively harnesses AI technologies to enhance defensive capabilities, it must also fully recognize the potential risks of autonomous agent technologies being misused, and devote efforts to developing detection and defence mechanisms specifically targeting autonomous agent-driven attacks.